# Sim-to-Real Transfer via 3D Feature Fields for Vision-and-Language Navigation

### v2

@@ 2 Related Works
Vision-and-Language Navigation. Vision and Language Navigation (VLN) [1, 2, 4, 19, 20, 11, 10] has gained significant attention in recent years. Early research focused on discrete environments due to the computational demands of exploring large action spaces in continuous environments. In discrete environments like the Matterport3D simulator [5], a predefined navigation connectivity graph is used. Agents observe panoramic RGB and depth images and can teleport to nearby nodes in the graph, navigating from a starting node to a target node following natural language instructions. Despite efficient training and high navigation performance, these settings are impractical for real-world applications. To address this, the continuous environment setting (VLN-CE [4]) was introduced. In VLN-CE, agents freely navigate any unobstructed location in 3D environments using low-level actions (e.g., turn left 15 degrees, turn right 15 degrees, or move forward 0.25 meters) in the Habitat simulator [6]. To overcome inefficiencies and poor performance in atomic action prediction, many works have designed subgoal or waypoint predictors to generate navigable waypoints, making VLN-CE more similar to discrete environments. By predicting several candidate waypoints, the agent selects the optimal one and executes atomic actions to move. This strategy significantly improves VLN-CE performance. However, these methods often rely on panoramic RGB-D images [15, 21, 14], limiting their applicability to real-world robots. In contrast, our work utilizes only a monocular RGB-D camera to construct a dynamic global semantic and occupancy map and then predicts the candidate waypoints, with better practicality.
@@
Sim-to-Real Transfer for VLN. Anderson [15] makes the first attempt at VLN sim-to-real transfer with panoramic cameras, which propose a subgoal model based on a 2D laser scanner to identify nearby waypoints. He finds that sim-to-real transfer to an unseen environment is challenging with no prior navigation graph. The core issue is that the navigable candidate waypoints predicted by the subgoal model exhibit significant deviations compared to the pre-defined navigation graph. More recently, to leverage the powerful reasoning and generalization capabilities of large language models or multimodal models to assist in real-world VLN, VLMaps [22] and DiscussNav [23] achieve VLN by combining multiple large models or foundation models. However, the navigation performance of these zero-shot methods is very limited. Therefore, NaVid [7] attempts to fine-tune the large model [24] using a large amount of video-based VLN data to achieve superior performance for real-world VLN with only a monocular camera. The large models for step-by-step atomic action prediction incur significant computational costs and response delays. Our work aims to design a universal solution to assist mainstream VLN models in sim-to-real transfer with practicality, high performance, and low cost.
@@
Generalizable 3D Feature Fields. The neural radiance field (NeRF) [17, 25] has gained significant popularity in various AI tasks, which predicts an image from an arbitrary viewpoint in a scene. However, the traditional NeRF methods with implicit MLP networks can only synthesize novel view images in seen scenes, which makes it difficult to generalize to unseen scenes and adapt to many embodied AI tasks. To avoid the difficulties of pixel-wise RGB reconstruction in unseen environments, GeFF [26] and HNR [12] attempt to encode 2D visual observations into 3D representations (called 3D feature fields) via the depth map. Using volume rendering [17], these models decode novel view representations from the 3D feature field and align them with open-world features (e.g., CLIP embeddings [18]). The 3D feature fields can generalize to unseen scenes, enabling real-time construction and dynamic updates, which assist with various embodied tasks. In this work, we utilize hierarchical feature fields of the HNR model [12] to provide panoramic perception capabilities for monocular robots.
@@ 3.2 Semantic Traversable Map for Waypoint Prediction
$\tau_{n}$ represents volume transmittance, and $\Delta_{n}$ is the distance between adjacent sampled points.
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grants 62125207, 62102400, 62272436, and U23B2012, in part by the National Postdoctoral Program for Innovative Talents under Grant BX20200338.

### v3

@@ 2 Related Works
Vision-and-Language Navigation. Vision and Language Navigation (VLN) [1, 2, 4, 19, 20, 21, 10, 22, 23] has gained significant attention in recent years. Unlike object goal navigation [24, 25, 26, 27, 28, 29], VLN agents must understand more complex natural language instructions and navigate to the described destination. Early VLN research focused on discrete environments due to the computational demands of exploring large action spaces in continuous environments. In discrete environments like the Matterport3D simulator [5], a predefined navigation connectivity graph is used. Agents observe panoramic RGB and depth images and can teleport to nearby nodes in the graph, navigating from a starting node to a target node following natural language instructions. Despite efficient training and high navigation performance, these settings are impractical for real-world applications. To address this, the continuous environment setting (VLN-CE [4]) was introduced. In VLN-CE, agents freely navigate any unobstructed location in 3D environments using low-level actions (e.g., turn left 15 degrees, turn right 15 degrees, or move forward 0.25 meters) in the Habitat simulator [6]. To overcome inefficiencies and poor performance in atomic action prediction, many works have designed subgoal or waypoint predictors to generate navigable waypoints, making VLN-CE more similar to discrete environments. By predicting several candidate waypoints, the agent selects the optimal one and executes atomic actions to move. This strategy significantly improves VLN-CE performance. However, these methods often rely on panoramic RGB-D images [15, 30, 14], limiting their applicability to real-world robots. In contrast, our work utilizes only a monocular RGB-D camera to construct a dynamic global semantic and occupancy map and then predicts the candidate waypoints, with better practicality.
@@
Sim-to-Real Transfer for VLN. Anderson [15] makes the first attempt at VLN sim-to-real transfer with panoramic cameras, which propose a subgoal model based on a 2D laser scanner to identify nearby waypoints. He finds that sim-to-real transfer to an unseen environment is challenging with no prior navigation graph. The core issue is that the navigable candidate waypoints predicted by the subgoal model exhibit significant deviations compared to the pre-defined navigation graph. More recently, to leverage the powerful reasoning and generalization capabilities of large language models or multimodal models to assist in real-world VLN, VLMaps [31] and DiscussNav [32] achieve VLN by combining multiple large models or foundation models. However, the navigation performance of these zero-shot methods is very limited. Therefore, NaVid [7] attempts to fine-tune the large model [33] using a large amount of video-based VLN data to achieve superior performance for real-world VLN with only a monocular camera. The large models for step-by-step atomic action prediction incur significant computational costs and response delays. Our work aims to design a universal VLN framework for sim-to-real transfer with practicality, high performance, and low cost.
@@
Generalizable 3D Feature Fields. The neural radiance field (NeRF) [17, 34] has gained significant popularity in various AI tasks, which predicts an image from an arbitrary viewpoint in a scene. However, the traditional NeRF methods with implicit MLP networks can only synthesize novel view images in seen scenes, which makes it difficult to generalize to unseen scenes and adapt to many embodied AI tasks. To avoid the difficulties of pixel-wise RGB reconstruction in unseen environments, GeFF [35] and HNR [12] attempt to encode 2D visual observations into 3D representations (called 3D feature fields) via the depth map. Using volume rendering [17], these models decode novel view representations from the 3D feature field and align them with open-world features (e.g., CLIP embeddings [18]). The 3D feature fields can generalize to unseen scenes, enabling real-time construction and dynamic updates, which assist with various embodied tasks. In this work, we utilize the HNR model [12] to provide panoramic perception capabilities for monocular robots.
@@ 3.2 Semantic Traversable Map for Waypoint Prediction
$\tau_{n}$ represents volume transmittance, and $\Delta_{n}$ is the distance between adjacent sampled points. $\textbf{R}_{(u,v)}$ represent the region feature at the $u$-th row and $v$-th column of the predicted feature map R.
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grants 62125207, 62102400, 62272436, and U23B2012, in part by Beijing Natural Science Foundation under Grant L242020, in part by the National Postdoctoral Program for Innovative Talents under Grant BX20200338.